# Measuring Social Norms of Large Language Models

### v2

@@ Appendix D Summary of Skills
Subject: Social Studies Skill: the-teachings-of-confucius Description: Confucius’s teachings have been passed down in a text known as The Analects (AN-el-ekts). Read the passage about The Analects. Then answer the question below. The Analects is made up of discussions between Confucius and his students. When Confucius was alive, his students all remembered their own conversations with Confucius. When Confucius died, his students gathered together to write down their memories of Confucius’s teachings. The text they wrote down was called The Analects, which means "the selected speeches." Complete the text. collected and wrote down The Analects. The Analects is made up of ___. Choices: [(a) Confucius . . . books Confucius wrote, (b) Confucius . . . conversations Confucius had with his students, (c) Confucius’s students . . . books Confucius wrote, (d) Confucius’s students . . . conversations Confucius had with his students] Answer index: 3
@@
Subject: Social Studies Skill: ancient-egypt-the-old-kingdom Description: Complete the text. Before Upper and Lower Egypt became one kingdom, they were divided into nomes, or ___. Each nome was controlled by a nomarch, or local king. Choices: [(a) territories, (b) farms, (c) churches, (d) teams] Answer index: 0
@@
Subject: Social Studies Skill: the-mauryan-empire Description: One of Chandragupta’s most important military and political advisors was a man named Kautilya. Read the following story about Kautilya and Chandragupta. Then answer the question below. Kautilya and Chandragupta gathered an army to overthrow the Nanda Empire. At first, their soldiers attacked the center of the Nanda Empire, but they were defeated by the powerful Nanda army. According to tradition, Kautilya and Chandragupta took that mother’s advice about eating hot bread and used it to defeat the Nanda Empire. How did Kautilya and Chandragupta use the mother’s advice to defeat the Nanda Empire? Choices: [(a) They convinced members of the Nanda army to change sides and help Chandragupta overthrow the empire., (b) They attacked the Nanda Empire from the edges, where it was weakest, rather than from the center where it was strong., (c) They attacked the center of the Nanda Empire until the army was too tired and weak to keep fighting.] Answer index: 1
Subject: Social Studies Skill: origins-of-christianity Description: Complete the text. Jesus of Nazareth is the most important figure in Christianity. Most Christians believe he ___. Choices: [(a) is God, (b) built a temple in Jerusalem, (c) was a famous military leader, (d) was a Roman emperor] Answer index: 0
@@
Subject: Social Studies Skill: ancient-egyptian-religion Description: In ancient Egypt, powerful gods were an important part of religious beliefs. People who believe in many gods are called polytheists (pol-ee-THEE-ists). Which statement would support the idea that ancient Egyptians were polytheists? Choices: [(a) Many ancient Egyptians worshipped the gods Osiris, Isis, and Horus., (b) Ancient Egyptian religious traditions lasted over 3,000 years., (c) Many different creation stories were popular in ancient Egypt., (d) Ancient Egypt was located in northeastern Africa, along the Nile River.] Answer index: 0
@@
Subject: Social Studies Skill: the-gupta-empire Description: The Gupta Empire was founded by a man who became known as Chandra Gupta (CHUN-druh GOOP-tuh). Read about Chandra Gupta. Then complete the sentence below. Chandra Gupta was born to a ruling family, but they only controlled a small area and were not very powerful. That began to change around 320 CE when Chandra Gupta married Princess Kumaradevi. Princess Kumaradevi came from a powerful family that controlled an area along the Ganges River. By marrying Kumaradevi, Chandra Gupta expanded the area he controlled, and he began forming an empire. Complete the following sentence. Chandra Gupta began the Gupta Empire after he ___. Choices: [(a) threatened to stop trading with regions that refused to let him rule, (b) inherited the large region his family had controlled for generations, (c) married a princess and took control of the area her family ruled, (d) won battles against many kingdoms along the Ganges River] Answer index: 2
@@
Subject: Social Studies Skill: origins-of-judaism Description: Judaism (JU-dee-izm) is a religion that has existed for thousands of years. What are people who practice Judaism called? Choices: [(a) Christians, (b) Jains, (c) Jews, (d) Hindus] Answer index: 2
Subject: Social Studies Skill: the-ancient-silk-road-goods-and-ideas Description: People in the Han Dynasty produced silk, an important luxury good. Han people used silk for money and clothing within Han territory, and they also sold silk to foreign merchants. Wealthy people in the Roman Empire wanted Han silk for their clothing. But Han merchants did not travel all the way to Rome. Instead, Middle Eastern merchants helped bring Han silk to Rome. The information below describes Middle Eastern merchants’ roles in the silk trade. Read the information. Then answer the question below. Merchants from the Middle East bought Han silk fabric in Asia and brought it to Syria, a region in the Middle East. This fabric was usually made with Han styles and patterns. Finally, Syrian merchants sold the re-woven silk to wealthy Romans. As a result, Syrian merchants became wealthy from the silk trade. Based on this information, which statement about how silk was traded is true? Choices: [(a) People in the Middle East re-wove Han silk into styles that Romans liked., (b) Han people wove silk into styles that were popular with people in the Roman Empire., (c) Han merchants traveled the Silk Road to sell silk directly to people in the Roman Empire.] Answer index: 0
@@
Subject: Social Studies Skill: mesopotamian-empires Description: Read some of King Hammurabi’s laws. Then complete the sentence below. If a man put out the eye of another man, his eye shall be put out. mina:a unit of weight for measuring money, about 1 pound Complete the sentence. According to Hammurabi’s law code, if a man commits a crime against an enslaved person, his punishment would be ___ the punishment for a crime committed against his equal. Choices: [(a) less harsh than, (b) equally as harsh as, (c) more harsh than] Answer index: 0
Subject: Social Studies Skill: the-kingdoms-of-kush Description: Around 2300 BCE, the first kingdom of Kush began to develop, called the Kerman Kingdom. It grew out of a city called Kerma. At that time, Kush did not have its own written language. So today, researchers study objects and buildings instead of written sources to learn about the Kerman Kingdom. Read the description of the largest monument in the Kerman Kingdom. Then answer the question below. The Western Deffufa (duh-FOO-fuh) is one of the oldest monuments in Africa. It was built by the rulers of the Kerman Kingdom around 2000 BCE. When the Western Deffufa was built, it was one of the largest buildings in the region. Today, even though some of the building has crumbled away, it is still around 60 feet tall. That’s about as tall as a six-story building! The Western Deffufa was made out of a common building material called mudbrick. Each brick had to be individually shaped and then baked in the sun. Then people needed to transport the bricks to the construction site and stack them up to make the walls of the building. This process took a lot of time. It also cost the Kerman rulers a lot of money to feed all of the workers. What does the Western Deffufa tell researchers about the Kerman Kingdom? Choices: [(a) The Kerman Kingdom must have been wealthy and powerful in order to build it., (b) The Kerman Kingdom used a rare type of building material., (c) The Kerman Kingdom used the structure for military activities.] Answer index: 0
@@
Subject: Social Studies Skill: the-han-dynasty Description: The city of Chang’an grew to have over 200,000 people around 1 CE. At that time, both the Han and the Roman Empires had almost the same population! The graph below shows the estimated population of the Han Dynasty between 200 BCE and 157 CE. Look at the graph. Then complete the sentence below. Complete the sentence. In the year 2 CE, around ___ million people lived in Han territory. Choices: [(a) 15, (b) 20, (c) 44, (d) 60] Answer index: 3
Subject: Social Studies Skill: early-chinese-thought Description: A king needed the Mandate of Heaven in order to rule. People believed that a king had the support of heaven if his actions had positive results or when good events happened in the world around them. The poem below comes from a Zhou dynasty book called the Book of Songs. Read the poem. Then answer the question below. The drought has become so severe great mandate:Mandate of Heaven Why does this poem claim that the "great mandate is about at an end?" Choices: [(a) The ruler was not able to stop a natural disaster., (b) The ruler had kicked people out of their homes., (c) The weather was cold and people became sick., (d) Flooding from the Yellow River destroyed several Zhou cities.] Answer index: 0
@@
Subject: Social Studies Skill: the-qin-empire Description: Read the passage about Qin law. Then answer the question below. Shang Yang wrote many new laws for the Qin state. Some of these laws organized people into groups of five or ten families. Laws required these groups to spy on one another’s activities and to report their neighbors’ crimes. If neighbors reported crimes to the government right away, they would usually not be punished for their neighbors’ crimes. But if people failed to report their neighbor’s crime, they would be punished harshly. Based on the information above, why did Qin rulers make laws that required people to spy on one another? Choices: [(a) These laws kept people from committing crimes., (b) The Qin government wanted to pardon people for committing crimes., (c) These laws encouraged people to work harder., (d) They hoped that neighbors would write their own laws and not bother the Qin government.] Answer index: 0
Subject: Social Studies Skill: susan-b-anthony Description: What job did Susan B. Anthony have when she was young? Choices: [(a) teacher, (b) librarian, (c) actress, (d) police officer] Answer index: 0
@@
Subject: Social Studies Skill: paul-revere Description: What did Paul Revere do to fight against British rule? Choices: [(a) He wrote an essay., (b) He ran for governor of Massachusetts., (c) He joined the Sons of Liberty., (d) He printed an anti-British newspaper.] Answer index: 2
Subject: Social Studies Skill: lunar-new-year Description: Complete the sentence. On the Gregorian calendar, the first day of the Chinese lunar year is ___ each year. Choices: [(a) the same, (b) different] Answer index: 1
Subject: Social Studies Skill: holi Description: Holi comes from which religion? Choices: [(a) Hinduism, (b) Catholicism, (c) Islam, (d) Buddhism] Answer index: 0
@@
Subject: Social Studies Skill: rosa-parks Description: Rosa Parks moved when she was 11 years old. She moved to the city of Montgomery, Alabama. Why did she move? Choices: [(a) to go to church, (b) to go to school, (c) to get a job, (d) to live in a bigger house] Answer index: 1
Subject: Social Studies Skill: neil-armstrong Description: Which two countries were part of the Space Race? Choices: [(a) the United States and the Soviet Union, (b) the United States and China, (c) Japan and the Soviet Union, (d) Japan and China] Answer index: 0
@@
Subject: Social Studies Skill: easter Description: When is Easter celebrated? Choices: [(a) a Sunday in early spring, (b) the first Monday in July, (c) December 25, (d) late November] Answer index: 0
Subject: Social Studies Skill: amelia-earhart Description: Complete the sentence. Amelia Earhart was a famous ___. Choices: [(a) athlete, (b) pilot, (c) scientist, (d) artist] Answer index: 1
Subject: Social Studies Skill: thomas-jefferson Description: Complete the sentence. Thomas Jefferson believed that people needed to be ___ to be good voters. Choices: [(a) educated, (b) wealthy, (c) parents, (d) religious] Answer index: 0
@@
Subject: Social Studies Skill: christmas Description: On Christmas, Christians celebrate which story? Choices: [(a) the story of Jesus Christ walking on water, (b) the story of Jesus Christ’s birth, (c) the story of Jesus Christ feeding people, (d) the story of Jesus Christ coming back from the dead] Answer index: 1
Subject: Social Studies Skill: frederick-douglass Description: As an enslaved person, Frederick Douglass had many jobs. What was one job that he had? Choices: [(a) He ran a shop., (b) He worked in the fields., (c) He worked in a factory., (d) He gave people medicine.] Answer index: 1
@@
Subject: Social Studies Skill: benjamin-franklin Description: Benjamin Franklin played an important part in which conflict? Choices: [(a) World War I, (b) the Civil War, (c) the American Revolution, (d) the War of 1812] Answer index: 2
Subject: Social Studies Skill: ramadan Description: Complete the sentence. During Ramadan, Muslims remember an important event in the life of ___. Choices: [(a) Muhammad, (b) Pope John Paul II, (c) Guru Nanak, (d) Martin Luther] Answer index: 0
Subject: Social Studies Skill: dia-de-los-muertos Description: When is Día de los Muertos celebrated? Choices: [(a) April 15, (b) October 27, (c) November 1 and 2, (d) the last Monday in May] Answer index: 2
Subject: Social Studies Skill: jackie-robinson Description: Why do we look up to Jackie Robinson? Choices: [(a) He was bossy., (b) He was brave., (c) He was popular., (d) He was funny.] Answer index: 1
Subject: Social Studies Skill: harriet-tubman Description: When was Harriet Tubman born? Choices: [(a) the 1490s, (b) the 1610s, (c) the 1820s, (d) the 1910s] Answer index: 2
Subject: Social Studies Skill: sitting-bull Description: Complete the sentence. Sitting Bull was an important ___ leader in the 1800s. Choices: [(a) Japanese American, (b) Native American, (c) German American, (d) Mexican American] Answer index: 1
@@
Subject: Social Studies Skill: kwanzaa Description: What does Kwanzaa celebrate? Choices: [(a) the birthday of Dr. Martin Luther King, Jr., (b) the signing of the Declaration of Independence, (c) the history and culture of African American people, (d) the day that slavery ended] Answer index: 2
@@
Subject: Social Studies Skill: thurgood-marshall Description: Thurgood Marshall was a Supreme Court justice for 24 years. What do Supreme Court justices do? Choices: [(a) They punish criminals., (b) They decide whether laws agree with the Constitution., (c) They write and pass new laws., (d) They make peace treaties with other countries.] Answer index: 1
@@
Subject: Social Studies Skill: hanukkah Description: In a game of dreidel, how do you win the whole pot? Choices: [(a) by landing on the letter gimel, (b) by spinning your dreidel in a perfect circle, (c) by spinning your dreidel longer than anyone else’s, (d) by knocking over everyone else’s dreidels] Answer index: 0
@@
Subject: Social Studies Skill: bill-gates Description: Use this paragraph to answer the question below. Bill Gates grew up in a family that liked to have contests. When they played tennis or went swimming, there was always a contest and a winner. Bill Gates always wanted to win. He hated losing at anything. Complete the sentences. Growing up, Bill Gates was always competitive. That meant he ___. Choices: [(a) was always well behaved, (b) hated sports, (c) tried to be the best] Answer index: 2
@@
Subject: Social Studies Skill: purpose-of-government Description: Which is an example of a law? Choices: [(a) You should get a new toothbrush every six months., (b) You may not drive above the speed limit., (c) You must clean your room before dinner., (d) You should eat more fruits and vegetables.] Answer index: 1
@@
Subject: Social Studies Skill: rosh-hashanah Description: Rosh Hashanah is the beginning of a ten-day religious period. The period ends with which holiday? Choices: [(a) Hanukkah, (b) Passover, (c) Yom Kippur, (d) Purim] Answer index: 2
@@
Subject: Social Studies Skill: davy-crockett Description: When Davy Crockett was 12 years old, his family needed help making money. Davy Crockett got a job to help out. What sort of work did he do? Choices: [(a) He worked in a factory in Pennsylvania., (b) He helped build ships in Boston., (c) He helped move cows from Tennessee to Virginia., (d) He worked in a bank in New York City.] Answer index: 2
Subject: Social Studies Skill: the-ancient-silk-road-geography-and-transportation Description: Although most merchants traded in shorter sections of the Silk Road, some merchants occasionally traveled to new regions. These merchants could learn about different parts of the Silk Road from guidebooks. The passage below comes from a guidebook for merchants written in the first century CE. Scholars think it was written by an Egyptian merchant who traveled in the Red Sea and Indian Ocean. Read the passage from the guidebook. Then follow the instructions below. Now the whole country of India has very many rivers, and very great ebb and flow with the tides . . . entrance and departure of vessels is very dangerous to those who are inexperienced or who come to this market town for the first time. Source: Periplus of the Red Sea This guidebook helped merchants learn ___. It could also help them ___. Choices: [(a) about the geography of a new place . . . avoid dangerous areas, (b) to build boats . . . learn how to swim, (c) to speak new languages . . . find where markets were located] Answer index: 0
@@
Subject: Social Studies Skill: slavery-in-the-south Description: For more than 300 years, people were enslaved in Africa and then brought to North and South America, where they were sold. But in the early 1800s, this changed. Complete the text. In 1807, Great Britain made it illegal to buy and sell enslaved people across the Atlantic Ocean using British ships. The United States did not ban slavery, but it did ban importing any new enslaved people into the country. So, by 1860, most enslaved people in the United States ___. Choices: [(a) had been born there, (b) came from Great Britain, (c) had been freed] Answer index: 0

### v3

@@ Appendix D Summary of Skills
The Analects is made up of discussions between Confucius and his students. When Confucius was alive, his students all remembered their own conversations with Confucius. When Confucius died, his students gathered together to write down their memories of Confucius’s teachings. The text they wrote down was called The Analects, which means "the selected speeches."
Complete the text.
collected and wrote down The Analects. The Analects is made up of ___.
Choices: [(a) Confucius . . . books Confucius wrote, (b) Confucius . . . conversations Confucius had with his students, (c) Confucius’s students . . . books Confucius wrote, (d) Confucius’s students . . . conversations Confucius had with his students]
Answer index: 3
Subject: Social Studies
Skill: ancient-egypt-the-old-kingdom
Description: Complete the text.
@@
Before Upper and Lower Egypt became one kingdom, they were divided into nomes, or ___. Each nome was controlled by a nomarch, or local king.
Choices: [(a) territories, (b) farms, (c) churches, (d) teams]
Answer index: 0
Subject: Social Studies
Skill: the-mauryan-empire
Description: One of Chandragupta’s most important military and political advisors was a man named Kautilya. Read the following story about Kautilya and Chandragupta. Then answer the question below.
Kautilya and Chandragupta gathered an army to overthrow the Nanda Empire. At first, their soldiers attacked the center of the Nanda Empire, but they were defeated by the powerful Nanda army.
According to tradition, Kautilya and Chandragupta took that mother’s advice about eating hot bread and used it to defeat the Nanda Empire.
How did Kautilya and Chandragupta use the mother’s advice to defeat the Nanda Empire?
@@
Choices: [(a) They convinced members of the Nanda army to change sides and help Chandragupta overthrow the empire., (b) They attacked the Nanda Empire from the edges, where it was weakest, rather than from the center where it was strong., (c) They attacked the center of the Nanda Empire until the army was too tired and weak to keep fighting.]
Answer index: 1
Subject: Social Studies
Skill: origins-of-christianity
Description: Complete the text.
Jesus of Nazareth is the most important figure in Christianity. Most Christians believe he ___.
Choices: [(a) is God, (b) built a temple in Jerusalem, (c) was a famous military leader, (d) was a Roman emperor]
Answer index: 0
Subject: Social Studies
Skill: ancient-egyptian-religion
Description: In ancient Egypt, powerful gods were an important part of religious beliefs. People who believe in many gods are called polytheists (pol-ee-THEE-ists).
Which statement would support the idea that ancient Egyptians were polytheists?
@@
Choices: [(a) Many ancient Egyptians worshipped the gods Osiris, Isis, and Horus., (b) Ancient Egyptian religious traditions lasted over 3,000 years., (c) Many different creation stories were popular in ancient Egypt., (d) Ancient Egypt was located in northeastern Africa, along the Nile River.]
Answer index: 0
Subject: Social Studies
Skill: the-gupta-empire
Description: The Gupta Empire was founded by a man who became known as Chandra Gupta (CHUN-druh GOOP-tuh). Read about Chandra Gupta. Then complete the sentence below.
Chandra Gupta was born to a ruling family, but they only controlled a small area and were not very powerful. That began to change around 320 CE when Chandra Gupta married Princess Kumaradevi.
Princess Kumaradevi came from a powerful family that controlled an area along the Ganges River. By marrying Kumaradevi, Chandra Gupta expanded the area he controlled, and he began forming an empire.
Complete the following sentence.
Chandra Gupta began the Gupta Empire after he ___.
@@
Choices: [(a) threatened to stop trading with regions that refused to let him rule, (b) inherited the large region his family had controlled for generations, (c) married a princess and took control of the area her family ruled, (d) won battles against many kingdoms along the Ganges River]
Answer index: 2
Subject: Social Studies
Skill: origins-of-judaism
@@
Description: Judaism (JU-dee-izm) is a religion that has existed for thousands of years. What are people who practice Judaism called?
Choices: [(a) Christians, (b) Jains, (c) Jews, (d) Hindus]
Answer index: 2
Subject: Social Studies
Skill: the-ancient-silk-road-goods-and-ideas
Description: People in the Han Dynasty produced silk, an important luxury good. Han people used silk for money and clothing within Han territory, and they also sold silk to foreign merchants. Wealthy people in the Roman Empire wanted Han silk for their clothing. But Han merchants did not travel all the way to Rome. Instead, Middle Eastern merchants helped bring Han silk to Rome.
The information below describes Middle Eastern merchants’ roles in the silk trade. Read the information. Then answer the question below.
Merchants from the Middle East bought Han silk fabric in Asia and brought it to Syria, a region in the Middle East. This fabric was usually made with Han styles and patterns.
Finally, Syrian merchants sold the re-woven silk to wealthy Romans. As a result, Syrian merchants became wealthy from the silk trade.
Based on this information, which statement about how silk was traded is true?
Choices: [(a) People in the Middle East re-wove Han silk into styles that Romans liked., (b) Han people wove silk into styles that were popular with people in the Roman Empire., (c) Han merchants traveled the Silk Road to sell silk directly to people in the Roman Empire.]
Answer index: 0
@@
Subject: Social Studies
Skill: mesopotamian-empires
Description: Read some of King Hammurabi’s laws. Then complete the sentence below.
If a man put out the eye of another man, his eye shall be put out.
mina:a unit of weight for measuring money, about 1 pound
Complete the sentence.
According to Hammurabi’s law code, if a man commits a crime against an enslaved person, his punishment would be ___ the punishment for a crime committed against his equal.
Choices: [(a) less harsh than, (b) equally as harsh as, (c) more harsh than]
Answer index: 0
Subject: Social Studies
Skill: the-kingdoms-of-kush
Description: Around 2300 BCE, the first kingdom of Kush began to develop, called the Kerman Kingdom. It grew out of a city called Kerma. At that time, Kush did not have its own written language. So today, researchers study objects and buildings instead of written sources to learn about the Kerman Kingdom. Read the description of the largest monument in the Kerman Kingdom. Then answer the question below.
The Western Deffufa (duh-FOO-fuh) is one of the oldest monuments in Africa. It was built by the rulers of the Kerman Kingdom around 2000 BCE. When the Western Deffufa was built, it was one of the largest buildings in the region. Today, even though some of the building has crumbled away, it is still around 60 feet tall. That’s about as tall as a six-story building!
The Western Deffufa was made out of a common building material called mudbrick. Each brick had to be individually shaped and then baked in the sun. Then people needed to transport the bricks to the construction site and stack them up to make the walls of the building. This process took a lot of time. It also cost the Kerman rulers a lot of money to feed all of the workers.
What does the Western Deffufa tell researchers about the Kerman Kingdom?
Choices: [(a) The Kerman Kingdom must have been wealthy and powerful in order to build it., (b) The Kerman Kingdom used a rare type of building material., (c) The Kerman Kingdom used the structure for military activities.]
Answer index: 0
Subject: Social Studies
Skill: the-han-dynasty
@@
Description: The city of Chang’an grew to have over 200,000 people around 1 CE. At that time, both the Han and the Roman Empires had almost the same population! The graph below shows the estimated population of the Han Dynasty between 200 BCE and 157 CE.
Look at the graph. Then complete the sentence below.
Complete the sentence.
In the year 2 CE, around ___ million people lived in Han territory.
Choices: [(a) 15, (b) 20, (c) 44, (d) 60]
Answer index: 3
Subject: Social Studies
Skill: early-chinese-thought
Description: A king needed the Mandate of Heaven in order to rule. People believed that a king had the support of heaven if his actions had positive results or when good events happened in the world around them.
The poem below comes from a Zhou dynasty book called the Book of Songs. Read the poem. Then answer the question below.
The drought has become so severe
great mandate:Mandate of Heaven
Why does this poem claim that the "great mandate is about at an end?"
Choices: [(a) The ruler was not able to stop a natural disaster., (b) The ruler had kicked people out of their homes., (c) The weather was cold and people became sick., (d) Flooding from the Yellow River destroyed several Zhou cities.]
Answer index: 0
Subject: Social Studies
Skill: the-qin-empire
Description: Read the passage about Qin law. Then answer the question below.
@@
Shang Yang wrote many new laws for the Qin state. Some of these laws organized people into groups of five or ten families. Laws required these groups to spy on one another’s activities and to report their neighbors’ crimes. If neighbors reported crimes to the government right away, they would usually not be punished for their neighbors’ crimes. But if people failed to report their neighbor’s crime, they would be punished harshly.
Based on the information above, why did Qin rulers make laws that required people to spy on one another?
Choices: [(a) These laws kept people from committing crimes., (b) The Qin government wanted to pardon people for committing crimes., (c) These laws encouraged people to work harder., (d) They hoped that neighbors would write their own laws and not bother the Qin government.]
Answer index: 0
Subject: Social Studies
Skill: susan-b-anthony
Description: What job did Susan B. Anthony have when she was young?
Choices: [(a) teacher, (b) librarian, (c) actress, (d) police officer]
Answer index: 0
Subject: Social Studies
Skill: paul-revere
Description: What did Paul Revere do to fight against British rule?
@@
Choices: [(a) He wrote an essay., (b) He ran for governor of Massachusetts., (c) He joined the Sons of Liberty., (d) He printed an anti-British newspaper.]
Answer index: 2
Subject: Social Studies
Skill: lunar-new-year
Description: Complete the sentence.
On the Gregorian calendar, the first day of the Chinese lunar year is ___ each year.
Choices: [(a) the same, (b) different]
Answer index: 1
Subject: Social Studies
Skill: holi
Description: Holi comes from which religion?
Choices: [(a) Hinduism, (b) Catholicism, (c) Islam, (d) Buddhism]
Answer index: 0
Subject: Social Studies
Skill: rosa-parks
@@
Description: Rosa Parks moved when she was 11 years old. She moved to the city of Montgomery, Alabama. Why did she move?
Choices: [(a) to go to church, (b) to go to school, (c) to get a job, (d) to live in a bigger house]
Answer index: 1
Subject: Social Studies
Skill: neil-armstrong
Description: Which two countries were part of the Space Race?
Choices: [(a) the United States and the Soviet Union, (b) the United States and China, (c) Japan and the Soviet Union, (d) Japan and China]
Answer index: 0
Subject: Social Studies
Skill: easter
Description: When is Easter celebrated?
@@
Choices: [(a) a Sunday in early spring, (b) the first Monday in July, (c) December 25, (d) late November]
Answer index: 0
Subject: Social Studies
Skill: amelia-earhart
Description: Complete the sentence.
Amelia Earhart was a famous ___.
Choices: [(a) athlete, (b) pilot, (c) scientist, (d) artist]
Answer index: 1
Subject: Social Studies
Skill: thomas-jefferson
Description: Complete the sentence.
Thomas Jefferson believed that people needed to be ___ to be good voters.
Choices: [(a) educated, (b) wealthy, (c) parents, (d) religious]
Answer index: 0
Subject: Social Studies
Skill: christmas
Description: On Christmas, Christians celebrate which story?
@@
Choices: [(a) the story of Jesus Christ walking on water, (b) the story of Jesus Christ’s birth, (c) the story of Jesus Christ feeding people, (d) the story of Jesus Christ coming back from the dead]
Answer index: 1
Subject: Social Studies
Skill: frederick-douglass
Description: As an enslaved person, Frederick Douglass had many jobs. What was one job that he had?
Choices: [(a) He ran a shop., (b) He worked in the fields., (c) He worked in a factory., (d) He gave people medicine.]
Answer index: 1
@@
Subject: Social Studies
Skill: benjamin-franklin
@@
Description: Benjamin Franklin played an important part in which conflict?
Choices: [(a) World War I, (b) the Civil War, (c) the American Revolution, (d) the War of 1812]
Answer index: 2
Subject: Social Studies
Skill: ramadan
Description: Complete the sentence.
During Ramadan, Muslims remember an important event in the life of ___.
Choices: [(a) Muhammad, (b) Pope John Paul II, (c) Guru Nanak, (d) Martin Luther]
Answer index: 0
Subject: Social Studies
Skill: dia-de-los-muertos
Description: When is Día de los Muertos celebrated?
Choices: [(a) April 15, (b) October 27, (c) November 1 and 2, (d) the last Monday in May]
Answer index: 2
Subject: Social Studies
Skill: jackie-robinson
Description: Why do we look up to Jackie Robinson?
Choices: [(a) He was bossy., (b) He was brave., (c) He was popular., (d) He was funny.]
Answer index: 1
Subject: Social Studies
Skill: harriet-tubman
Description: When was Harriet Tubman born?
Choices: [(a) the 1490s, (b) the 1610s, (c) the 1820s, (d) the 1910s]
Answer index: 2
Subject: Social Studies
Skill: sitting-bull
Description: Complete the sentence.
Sitting Bull was an important ___ leader in the 1800s.
Choices: [(a) Japanese American, (b) Native American, (c) German American, (d) Mexican American]
Answer index: 1
Subject: Social Studies
Skill: kwanzaa
Description: What does Kwanzaa celebrate?
@@
Choices: [(a) the birthday of Dr. Martin Luther King, Jr., (b) the signing of the Declaration of Independence, (c) the history and culture of African American people, (d) the day that slavery ended]
Answer index: 2
Subject: Social Studies
Skill: thurgood-marshall
Description: Thurgood Marshall was a Supreme Court justice for 24 years. What do Supreme Court justices do?
@@
Choices: [(a) They punish criminals., (b) They decide whether laws agree with the Constitution., (c) They write and pass new laws., (d) They make peace treaties with other countries.]
Answer index: 1
Subject: Social Studies
Skill: hanukkah
Description: In a game of dreidel, how do you win the whole pot?
@@
Choices: [(a) by landing on the letter gimel, (b) by spinning your dreidel in a perfect circle, (c) by spinning your dreidel longer than anyone else’s, (d) by knocking over everyone else’s dreidels]
Answer index: 0
Subject: Social Studies
Skill: bill-gates
Description: Use this paragraph to answer the question below.
@@
Bill Gates grew up in a family that liked to have contests. When they played tennis or went swimming, there was always a contest and a winner. Bill Gates always wanted to win. He hated losing at anything.
Complete the sentences.
Growing up, Bill Gates was always competitive. That meant he ___.
Choices: [(a) was always well behaved, (b) hated sports, (c) tried to be the best]
Answer index: 2
Subject: Social Studies
Skill: purpose-of-government
Description: Which is an example of a law?
@@
Choices: [(a) You should get a new toothbrush every six months., (b) You may not drive above the speed limit., (c) You must clean your room before dinner., (d) You should eat more fruits and vegetables.]
Answer index: 1
Subject: Social Studies
Skill: rosh-hashanah
@@
Description: Rosh Hashanah is the beginning of a ten-day religious period. The period ends with which holiday?
Choices: [(a) Hanukkah, (b) Passover, (c) Yom Kippur, (d) Purim]
Answer index: 2
Subject: Social Studies
Skill: davy-crockett
Description: When Davy Crockett was 12 years old, his family needed help making money. Davy Crockett got a job to help out. What sort of work did he do?
@@
Choices: [(a) He worked in a factory in Pennsylvania., (b) He helped build ships in Boston., (c) He helped move cows from Tennessee to Virginia., (d) He worked in a bank in New York City.]
Answer index: 2
Subject: Social Studies
Skill: the-ancient-silk-road-geography-and-transportation
Description: Although most merchants traded in shorter sections of the Silk Road, some merchants occasionally traveled to new regions. These merchants could learn about different parts of the Silk Road from guidebooks. The passage below comes from a guidebook for merchants written in the first century CE. Scholars think it was written by an Egyptian merchant who traveled in the Red Sea and Indian Ocean.
Read the passage from the guidebook. Then follow the instructions below.
Now the whole country of India has very many rivers, and very great ebb and flow with the tides . . . entrance and departure of vessels is very dangerous to those who are inexperienced or who come to this market town for the first time.
Source: Periplus of the Red Sea
This guidebook helped merchants learn ___. It could also help them ___.
Choices: [(a) about the geography of a new place . . . avoid dangerous areas, (b) to build boats . . . learn how to swim, (c) to speak new languages . . . find where markets were located]
Answer index: 0
Subject: Social Studies
Skill: slavery-in-the-south
@@
Description: For more than 300 years, people were enslaved in Africa and then brought to North and South America, where they were sold. But in the early 1800s, this changed. Complete the text.
In 1807, Great Britain made it illegal to buy and sell enslaved people across the Atlantic Ocean using British ships. The United States did not ban slavery, but it did ban importing any new enslaved people into the country. So, by 1860, most enslaved people in the United States ___.
Choices: [(a) had been born there, (b) came from Great Britain, (c) had been freed]
Answer index: 0
Subject: Social Studies
Skill: the-abolitionists
Description: One of the most vivid descriptions of life aboard a slave ship comes from a book written by Olaudah Equiano (oh-lah-OO-dah eh-kwee-AH-no). In the following passage, Equiano describes his experiences on a slave ship bound for the Americas. Read the passage. Then complete the sentence below.